OpenStudy (anonymous):

Solve each quadratic equation and write answers in simplest radical form. 3x^2+2=4x

OpenStudy (anonymous):

Move the "$x" to the LHS so that the quadratic equation is equal to zero. \[3x^2-4x+2=0\] Use the quadratic formula. \[x=\frac{-b\pm\sqrt{b^2-4ac}}{2a}\]

OpenStudy (anonymous):

"4x"*

OpenStudy (anonymous):

I get a negative 8 for the b^2-4(a)(c)

OpenStudy (anonymous):

If you got negative you would mark your answer as an imaginary number using i

OpenStudy (anonymous):

But my teacher said i have to find the greatest commonfactor of the negative and plug it on the formula for example in this problem o got 2 square root of 2 i'm just loat on what to do next

OpenStudy (anonymous):

You have to use the quadratic formula to solve. You can't factor it. So you plug it into the formula and get: \[4\pm \sqrt{16-(4)(3)(2)}/(2(3))\] \[4 \pm 2i \sqrt{2}/ 6\] then you simplify to get \[2 \pm i \sqrt{2}/ 3\] which looks like this if you want a drawing: